OpenStudy (mortonsalt):

Hello everyone! How do I find the horizontal and vertical asymptotes of each curve? (Will post equation next.)

OpenStudy (mortonsalt):

\[\frac{ x-9 }{ \sqrt{4x^2+3x+2} }\]

OpenStudy (campbell_st):

well I'd say it has no vertical asymptotes as the denominator will always be positive there are horizontal asymptotes something like y = -1/2 and y = 1/2 I just approximated the denominator to -2x and 2x I then uses y = x/2x and y = x/-2x

OpenStudy (mortonsalt):

Ohhh, alright! Thank you! :)